# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v4

@@ Abstract
Milgrom’s Modified Newtonian Dynamics (MOND) correction to Newtonian gravity or inertia is shown to be equivalent to a more fundamental transformation between a non-inertial local reference frame and the fixed background of the observable universe, complying with Mach’s principle. Both Newton’s gravitational constant and Milgrom’s MOND acceleration parameter or scale constant are substituted for two varying, measurable, physical, and cosmological parameters under the justification of Mach’s principle: causally connected mass and size of the universe. This Machian interpretation of MOND is mass, length, and time scale invariant at all regimes and free from fundamental constants and free parameters with the exception of the speed of light as the speed of causality and gravity. By extending the external field effect to include that of the rest of the universe, the correction is based on relative field intensities of the small and large scales. It respects the Machian effect by which, in absence of a background, rotational speed is undefined up to the speed of light. The Machian MOND interpretation is a necessary feature that a phenomenological non-linear theory of modified inertia or modified gravity which incorporates Mach’s principle, in agreement with galaxy rotation curves, should effectively reduce to as an approximation.
@@ 1.1 Modified Newtonian Dynamics MOND
The solution to the problem of the perihelion precession of Mercury first involved various failed attempts proposing the unobserved existence of a new small planet Vulcan, an asteroid belt, or a gas cloud orbiting the Sun. However, the resolution to the problem came as a modification to the laws of gravity by General Relativity (GR). Similarly, the observation of rotational velocities in disks of galaxies higher than predicted by Newtonian gravity, as depicted in galaxy rotation curves, has driven further research into modifications to the laws of gravity and mechanics. The discrepancies between observed luminance to gravitational mass ratio in galaxies and their missing Keplerian falloff in their velocity curves, together with faster radial velocities of galaxies in galaxy clusters than Newtonian predictions, among other evidence, are referred to as the ‘dark matter’ problem [1, 2, 3]. Modifying gravity or inertia instead of proposing the existence of physical cold dark matter is motivated by the Tully-Fisher law (the asymptotic velocities in rotation curves are proportional to the fourth root of their mass), the Renzo’s rule (features in visible mass strongly correlate with features in velocity profiles), and the cuspy halo problem (Newtonian gravity correctly predicts velocity profiles at galactic cores). These recent modifications to the classical law of gravity or laws of mechanics are called Modified (or Mondified) Newtonian Dynamics (MOND) models (or MilgrOmiaN Dynamics), with the first and main MOND being Milgrom’s MOND [4, 5, 6], which can be thought of as a modification to inertia (applied to Newton’s second law of motion) or to the inverse square law of gravity. Almost all full-fledged versions of MOND are just modified gravity, in which the inverse-square law of gravity approximately changes to an inverse linear law, and not modified inertia. As modified inertia, MOND modifies the kinetic term in the particle Lagrangian $mv^{2}/2$ or the ‘inertia term’ $mg$, and measured kinematic accelerations are not simply proportional to forces per unit mass.
@@
The reason for the anomalous motion of Mercury was that Newtonian gravity is not a good enough approximation for strong gravitational fields, such as near the Sun, where GR effects become significant. Milgrom’s MOND exceptionally explains the observed rotation curves of galaxies by imposing a correction based on two observational relationships: the Tully-Fisher law, and that the mass (or velocity) discrepancies are always observed below a particular acceleration scale. The simplest correction that satisfies these observational conditions results in a constant asymptotic rotational velocity $v=\sqrt[4]{GMa_{0}}$ with the need to introduce a constant of acceleration $a_{0}$. The connection between the classical Newtonian regime where $v=\sqrt{GM/r}$ and Milgrom’s deep MOND or low acceleration regime is determined by an ‘interpolating function’ $\mu\left(g/a_{0}\right)=1/(1+\left(a_{0}/g\right)^{n})^{(1/n)}$ (with $n=1$ for the simple and $n=2$ for the standard interpolating functions) based on $g$ as the ‘true acceleration’ and a constant of acceleration $a_{0}$ (also referred to as the Hubble acceleration or the acceleration scale constant), which is a free parameter adjusted by data fitting and is in principle considered a fundamental and universal constant in Milgrom’s MOND. The interpolating function must satisfy the conditions $\mu(x)\rightarrow 1$ for $x\ll 1$ to recover Newtonian laws and $\mu(x)\rightarrow x$ for $x\gg 1$ to satisfy the Tully-Fisher law. A fixed-distance or fixed-mass (independent of the system under study) modification alone cannot account for the observed rotation velocities due to the Tully-Fisher law, i.e., it is not possible to simply set a constant of distance or a constant of mass to modify the laws of gravity to fit all rotation curves of galaxies (because some galaxies exhibit the effect of dark matter at small radius while others at large radius, and same happens with different galaxy masses). However, combining both mass and distance in acceleration or gravitational field intensity, Milgrom’s MOND achieves this limit where the effect of dark matter in galaxies is generally observed. Still, Milgrom’s MOND, being insufficient to explain all dark matter related phenomena such as galaxy cluster dynamics, [7, 8], being non-relativistic, and not explaining phenomenologically the origin of the interpolating function, is an effective theory or approximation to a more fundamental modification to gravitation or inertia.
@@
The acceleration constant in Milgrom’s MOND $a_{0}\sim 10^{-10}\ m/s^{2}$, was already related to the cosmological scale through the Hubble constant $H_{0}$ (or the cosmological constant $\Lambda$) $a_{0}\sim cH_{0}\sim c^{2}\sqrt{\Lambda}\sim c^{2}/R_{u}$ originally by Milgrom [9], stating that this coincidence could point to a basic theory underlying MOND’s phenomenology. In particular, he stated that “an attractive possibility is that MOND results as a non-relativistic, small-scale expression of a fundamental theory by which inertia is a vestige of the interaction of a body with ‘the rest of the Universe’, in the spirit of Mach’s principle.”, and “We thus envisage inertia as resulting from the interaction of the accelerated body with some agent field, perhaps having to do with the vacuum fields, perhaps with an “inertia field” whose source is matter in the “rest of the universe”–in the spirit of Mach’s principle[10], although he did not further develop his theory in relation to Machian ideas.
@@
Progress in fundamental physics has been made by reducing the number of fundamental constants (such as the gravitational acceleration of the Earth and other planets for a universal gravitational constant, the unification of the speed of light with vacuum permittivity and permeability, and atomic constants for the Planck constant). Therefore, it is of considerable interest to explore modifications of the laws of gravity or mechanics without the need for more fundamental constants.
@@
Various theories of modified gravity or inertia are based on Milgrom’s MOND acceleration constant: John Moffat’s MOG scalar–tensor–vector gravity approach to galaxy rotation curves is based on two parameters, which are set to fit galactic Milgrom’s MOND’s acceleration constant [11], Erik’s Verlinde’s entropic gravity attempts to propose an underlying framework for Milgrom’s MOND’s acceleration constant [12], and Deur’s self-interacting gravitons model correction to Newtonian gravity contains a physical constant counterpart to Milgrom’s MOND’s acceleration constant in the form of $a=\Lambda/a_{0}=\sqrt{\Lambda}/c^{2}$ [13]. Others, such as quantized inertia, rely on the acceleration parameter $a_{0}=c^{2}/R_{u}$ where $R_{u}$ is the radius of the cosmological comoving horizon [14].
@@ 1.2 Mach’s Principle and modified inertia
The definition of absolute space, time, and motion suffer from circular reasoning. In Newtonian mechanics, absolute space is defined as the frame absent of inertial forces (also referred to as fictitious forces, which come from our choice of reference frame that is rotating in absolute space), but the absence of inertial forces is used to prove and identify the existence of absolute space in Newton’s rotating bucket of water experiment. According to Newton’s first law of motion, an object moves inertially if it is free from outside influences, but the fact that it is free from outside influences is inferred only by observing that it moves inertially. In Special Relativity (SR), light travels at the same speed in all inertial frames and absolute acceleration is defined relative to inertial frames, but inertial frames are defined as frames absent of acceleration (accelerometers always measure acceleration with respect to a reference frame of calibration, that must be assumed to be inertial). If a body is measured to have acceleration in one inertial frame, all other inertial frames will agree that it is accelerating.
@@
Leibniz’s relative space, time and motion ideas were further developed by Ernst Mach in his 1883 The Science of Mechanics, in which he criticized Newton’s conclusion of his bucket of water experiment stating that “No one is competent to say how the experiment would turn out if the sides of the vessel increased in thickness and mass till they were ultimately several leagues thick.[15]. According to Mach, the relative rotation of the water with respect to the bucket produces no noticeable centrifugal forces, and such forces are instead produced by its relative rotation with respect to Earth and the other celestial bodies (in this way, the relative motion between the several leagues thick bucket and the water could produce noticeable centrifugal forces in the water, and perhaps their non-relative motion could reduce these forces when both rotate with respect to the rest of the universe). In contrast to Newton, who attempted to explain the physical effects of inertia through a sort of resistance to motion within an unobservable absolute space with no physical properties, Mach conceived inertia as an interaction that required other external bodies to manifest. Newton did not suspect that the change in the water’s shape could be due to rotation relative to the rest of the universe because his action at a distance gravitational force from his infinite and homogeneous universe acting on the water cancels out according to Newton’s shell theorem and is independent of velocity or acceleration. But no local phenomena can ever be isolated from the rest of the universe, and its effect certainly reaches the water even if the sum of these forces is zero. According to Mach, and in opposition to Newton’s conclusions, a spinning bucket of water in an empty universe would not change the water’s shape (one could not detect relative motion in an empty universe, it would be undefinable), and if the rest of the universe was set spinning while the bucket was at rest, the water surface would curve, as both the spinning of the bucket and water or the spinning of the universe are indistinguishable systems without an absolute space. Thus, for Mach, the relativity symmetry of acceleration is broken by the existence of the background universe.
@@
Moreover, there is an exact coincidence between the local measurement of the angular velocity of the Earth through Foucault’s pendulum and the cosmological measurement through the apparent movement of distant stars and galaxies, which Newtonian gravity and mechanics cannot explain because it does not causally connect both measurements (and consider the determination of inertial frames by the fixed stars and Foucault’s pendulum a coincidence), but constitutes the basic idea of Mach’s principle: “The universe, as represented by the average motion of distant galaxies, does not appear to rotate relative to local inertial frames.[16].
@@
According to Mach, and known as Mach’s principle, the inertia of a body is not an independent and intrinsic property of matter (unlike in Newtonian mechanics, where a particle in an empty universe has the usual inertial properties), but rather the result of the action of the universe as a whole. Mach suggested that the fixed background distribution of matter in the universe must exert the inertial forces on a local accelerating body. In this way, the reference system with respect to which the universe is at rest or in uniform and rectilinear motion is a true inertial reference system. In Ernst Mach words, “I have remained to the present day the only one who insists upon referring the law of inertia to the Earth, and in the case of motions of great spatial and temporal extent to the fixed stars.[15]. Hence, inertial frames should be defined with respect to this rest frame, and local physical laws must be determined by the large-scale structure of the universe.
@@
Einstein’s pursuit of a relativistic gravitational theory was inspired by Mach, and he envisioned GR to fulfill his interpretation of Mach’s principle: “the metric tensor is completely caused and determined by the stress-energy tensor[17]. He noted that following Mach, it was expected that “1. The inertia of a body must increase when ponderable masses are piled up in its neighbourhood, 2. A body must experience an accelerating force when neighbouring masses are accelerated, and, in fact, the force must be in the same direction as the acceleration, 3. A rotating hollow body must generate inside of itself a “Coriolis field”, which deflects moving bodies in the sense of the rotation, and a radial centrifugal field as well[18]. He had hopes that by using tensors in GR, he would achieve the strong Machian version of relativity (in which the gravitational influence of the whole universe gave rise to inertia), but although points 2. and 3. were satisfied in GR (at least phenomenologically by frame-dragging), the first condition was not, because the immediate consequence of it is that a body in an otherwise empty universe should have no inertia, and the empty solution where the stress-energy tensor is zero everywhere in GR is the flat Minkowski spacetime of SR, in which test bodies have the usual inertia. Einstein then introduced the cosmological constant term (as the mass density o the universe) in his field equations in the hope that, with the cosmological term, they would have no solutions for a zero stress-energy tensor. However, de Sitter found in 1917 a solution for the field equations with the cosmological constant and with zero stress-energy tensor for an expanding universe, Einstein dismissed the cosmological constant as no longer justified, and he abandoned the ideas of Mach.
@@
Einstein’s equivalence principle (a frame linearly accelerated relative to an inertial frame in the Minkowskian spacetime of SR is locally identical to a frame at rest in a uniform gravitational field) axiomatically assumes that inertial and gravitational masses are equal. Gravity is the only infinite range force that cannot be screened due to the existence of only positive masses, unlike in electromagnetism, where positive and negative charges result in balance and neutrality, and astrophysical and cosmological structures are therefore not governed by electromagnetism. Moreover, there always exists a reference frame in which inertial forces vanish, just as for the case of gravitational forces in free fall. These equivalences indicate that both inertial and gravitational forces are of the same nature. Following Mach, in any two-body interaction, the influence of all other matter inside their causal spacetime should be taken into account. Inertia is thus a form of gravitational induction, appearing when a body is accelerated with respect to the rest of the universe, subjected to retarded action. But Newtonian gravity and GR were developed without considering today’s acknowledgment of the mass content and size of the observable universe. Then, it seems that to describe, for example, the Earth’s motion around the Sun, only local masses and distances are required, but the Machian perspective implies that the universe’s action is already taken into account in the Newtonian laws. The only way this can be true is through the Newtonian gravitational constant, which, together with the choice of the inertial frame of absolute space, both constitute the two arbitrary choices of Newton.
@@
In Newtonian gravity, the gravitational constant $G$ is the only fundamental constant of the theory, and it is known with far less precision than any other fundamental dimensional constant in physics. The uncertainty in its measurement, together with its problematic in Quantum Field Theory in the Planck energy (which is used as a cutoff for the energy density of the vacuum and the theoretical quantum corrections estimation of the Higg’s mass, leading to the vacuum catastrophe and hierarchy problems) has led to questioning its constant nature repeatedly over the last century; see [19] for constraints. GR carries this constant (a varying $G$ would imply a violation of the strong equivalence principle, but only the weak principle is supported by the very precise Eötvös experiments) with the addition of the speed of light, which can be derived from the vacuum permittivity and permeability constants. Therefore, it is reasonable to question whether the gravitational constant is also a derived parameter.
@@
The first historically suggestion of gravity arising necessarily as a consequence of the relativity of inertia came from Hans Reissner [20]. Schrödinger soon identified in 1925 [21] that GR did not fully implement Mach’s principle, and proposed a relationship between the gravitational potential of distant masses and the speed of light: “This remarkable relationship states that the (negative) potential of all masses at the point of observation, calculated with the gravitational constant valid at that observation point, must be equal to half the square of the speed of light.”. This conclusion was reached by imposing that the kinetic energy term had an origin in a potential-like interaction (such as all other forms of energy, which have an origin in an interaction) and was relational, so that $K=mv^{2}/2$ based on absolute mass and absolute velocity can be expressed through the gravitational Newtonian potential $Gm/r$ so that $K=(Gm_{i}/r_{ij})(m_{j}v_{ij}^{2}/2c^{2})$ with the need to introduce the square of the speed of light. The term $Gm_{i}/r_{ij}c^{2}=GM_{u}/c^{2}R_{u}\sim 1$ is considered missing in the original Newtonian formulation, with $M_{u}$ being the sum of all masses within $R_{u}$, the distance between the moving particle $m_{j}$ and $M_{u}$. Schrödinger’s formulation predicted a finite observable universe, since the potential of an infinite non-expanding eternal universe would also be infinite, even though the expansion of the universe was not known at the time. This relationship was later credited to Reissner’s 1915 work by Schrödinger himself.

### v5

@@ Abstract
Moderhai Milgrom’s Modified Newtonian Dynamics (MOND) correction to Newtonian gravity or inertia is shown to be equivalent to a more fundamental formulation considering a non-inertial local reference frame and the fixed background of the observable universe, in the spirit of Mach’s principle. Both Newton’s gravitational constant $G\sim c^{2}/(M_{u}/R_{u})$ and Milgrom’s MOND acceleration scale constant $a_{0}\sim GM_{u}/R_{u}^{2}$ are replaced by two varying, measurable, and cosmological quantities determined by the causally connected mass and size of the universe. They arise from an inverse and an inverse squared distance scalar fields of matter density, respectively. This Machian interpretation of MOND is invariant under global rescalings of mass, length, and time across all regimes and is free from fundamental constants and free parameters, except for the speed of light. Machian MOND satisfies the fundamental consequences of Mach’s principle not featured in Newton’s and Einstein’s theories: the decrease of inertia of a body when masses are removed from its neighborhood, and in the absence of a cosmic background, rotational motion is undefined up to the speed of light. Consequently, Machian MOND provides the necessary limiting behavior to which any phenomenological non-linear theory of modified inertia or gravity that incorporates Mach’s principle, in agreement with galaxy rotation curves, should reduce as an effective approximation.
@@ 1.1 Modified Newtonian Dynamics MOND
The solution to the problem of the perihelion precession of Mercury first involved various failed attempts proposing the unobserved existence of a new small planet Vulcan, an asteroid belt, or a gas cloud orbiting the Sun. However, the resolution came as a modification to the laws of gravity by General Relativity (GR). Similarly, the observation of rotational velocities in disks of galaxies higher than predicted by Newtonian gravity, as depicted in galaxy rotation curves, has driven further research into modifications to the laws of gravity and mechanics. The discrepancies between observed luminance to gravitational mass ratio in galaxies and their missing Keplerian falloff in their velocity curves, together with faster radial velocities of galaxies in galaxy clusters than Newtonian predictions, among other evidence, are referred to as the ‘dark matter’ problem [Zwicky1933, Rubin1983, Review]. Modifying gravity or inertia instead of proposing the existence of physical cold dark matter is motivated by the Tully-Fisher (TF) law (the asymptotic velocities in rotation curves are proportional to the fourth root of their baryonic mass), Renzo’s rule (features in visible mass strongly correlate with features in velocity profiles), and the cuspy halo problem (Newtonian gravity generally predicts velocity profiles in galactic cores with visible mass only). These recent modifications to the classical law of gravity or laws of mechanics are called Modified Newtonian Dynamics (MOND) models (or MilgrOmiaN Dynamics), with the first and main MOND being Milgrom’s MOND [Milgrom1983a, Milgrom1983b, Milgrom1983c], which can be thought of as a modification to inertia (applied to Newton’s second law of motion) or to the inverse square law of gravity. Almost all full-fledged versions of MOND are just modified gravity, in which the inverse-square law of gravity approximately changes to an inverse law, and not modified inertia. As modified inertia, MOND modifies the kinetic term in the particle Lagrangian $mv^{2}/2$ or the ‘inertia term’ $mg$, and measured kinematic accelerations are not simply proportional to forces per unit mass.
@@
The reason for the anomalous motion of Mercury was that Newtonian gravity is not a good enough approximation for strong gravitational fields, such as near the Sun, where non-linear GR effects become significant. Milgrom’s MOND exceptionally explains the observed rotation curves of galaxies by imposing a correction based on two observational relationships: the TF law, and that the mass or velocity discrepancies (between observed and Newtonianly estimated) are always observed below a particular acceleration scale. The simplest correction that satisfies these observational conditions results in an asymptotic rotational velocity $v_{\infty}=\sqrt[4]{GMa_{0}}$ independent of distance, with the need to introduce a constant of acceleration $a_{0}$. The connection between the classical Newtonian regime where $v=\sqrt{GM/r}$ and Milgrom’s deep MOND or low acceleration regime is determined by an ‘interpolating function’ $\mu\left(g/a_{0}\right)=1/(1+\left(a_{0}/g\right)^{n})^{(1/n)}$ (with $n=1$ for the simple and $n=2$ for the standard interpolating functions) based on $g$ as the ‘true acceleration’ and a constant of acceleration $a_{0}$ (also referred to as the Hubble acceleration or the acceleration scale constant), which is a free parameter adjusted by data fitting and is in principle considered a fundamental and universal constant in Milgrom’s MOND. At the deep MOND regime, $g=\sqrt{g_{N}a_{0}}$ with $g_{N}$ the Newtonian gravitational acceleration, and the effective potential is logarithmic at large distances $V(r)\sim\sqrt{GMa_{0}}\ln(r)$, diverging at infinity. Milgrom’s MOND is only length and time scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$ to achieve constant velocities independent of radius at the flat part of the galaxy rotation curves. The interpolating function must satisfy the conditions $\mu(x)\rightarrow 1$ for $x\gg 1$ to recover Newtonian laws and $\mu(x)\rightarrow x$ for $x\ll 1$ to satisfy the TF law. A fixed-distance or fixed-mass (independent of the system under study) modification alone cannot account for the observed rotation velocities due to the TF law, i.e., it is not possible to simply set a constant of distance or a constant of mass to modify the laws of gravity to fit all rotation curves of galaxies (because some galaxies exhibit the effect of dark matter at small radius while others at large radius, and the same happens with different galaxy masses). However, combining both mass and distance in acceleration or gravitational field intensity, Milgrom’s MOND achieves this limit where the effect of dark matter in galaxies is generally observed. Still, Milgrom’s MOND, being insufficient to explain all dark matter related phenomena such as galaxy cluster dynamics, [Sanders1998, Sanders2003], being non-relativistic, and not explaining phenomenologically the origin of the interpolating function, is an effective theory or approximation to a more fundamental modification to gravity or inertia.
@@
The acceleration constant in Milgrom’s MOND $a_{0}\sim 10^{-10}\ m/s^{2}$, was already related to the cosmological scale through the Hubble constant $H_{0}$ (or the cosmological constant $\Lambda$) $a_{0}\sim cH_{0}\sim c^{2}\sqrt{\Lambda}\sim c^{2}/R_{u}$ originally by Milgrom [Milgrom1983a], stating that this coincidence could point to a basic theory underlying MOND’s phenomenology. In particular, he stated that “an attractive possibility is that MOND results as a non-relativistic, small-scale expression of a fundamental theory by which inertia is a vestige of the interaction of a body with ‘the rest of the Universe’, in the spirit of Mach’s principle.”, “It seems to me that in looking for an ultimate theory, the Mach principle may serve as a most useful guide”, “The possible connection between such numerical “coincidences,” which relate parameters of “local” physics to cosmological parameters and the Mach principle, is quite obvious”, and “We thus envisage inertia as resulting from the interaction of the accelerated body with some agent field, perhaps having to do with the vacuum fields, perhaps with an “inertia field” whose source is matter in the “rest of the universe”–in the spirit of Mach’s principle[Milgrom1993, Milgrom1983a], although he did not further develop MOND in relation to Machian ideas.
@@
Progress in fundamental physics has been made by reducing the number of fundamental constants (such as the gravitational acceleration of the Earth and other planets for a universal gravitational constant, the unification of the speed of light with the vacuum permittivity and permeability, and atomic constants for the Planck constant). Therefore, it is of considerable interest to explore modifications of the laws of gravity or mechanics by removing fundamental constants.
@@
Various theories of modified gravity or inertia are based on Milgrom’s MOND acceleration constant: John Moffat’s MOG scalar–tensor–vector gravity approach to galaxy rotation curves is based on two parameters, which are set to fit the galactic Milgrom’s MOND’s acceleration constant [Brownstein_2006], Erik’s Verlinde’s entropic gravity attempts to propose an underlying framework for Milgrom’s MOND’s acceleration constant [Verlinde2017], and Deur’s self-interacting gravitons model correction to Newtonian gravity contains a physical constant counterpart to Milgrom’s MOND’s acceleration scale constant [Deur2003]. Others, such as quantized inertia, rely on the acceleration parameter $a_{0}=c^{2}/R_{u}$ where $R_{u}$ is the radius of the cosmological comoving horizon [McCulloch2007].
@@ 1.2 Mach’s Principle and modified inertia
The definition of absolute space, time, and motion suffer from circular reasoning. In Newtonian mechanics, absolute space is defined as the frame absent of inertial forces (also referred to as fictitious forces, which come from our choice of reference frame), but the absence of inertial forces is used to prove and identify the existence of absolute space in Newton’s rotating bucket of water experiment. According to Newton’s first law of motion, an object moves inertially if it is free from outside influences, but the fact that it is free from outside influences is inferred only by observing that it moves inertially. How can the inertial system be unequivocally identified? In Special Relativity (SR), light is measured to travel at the same speed in all inertial frames, and absolute acceleration is defined relative to inertial frames. But inertial frames are defined as frames absent of acceleration, and accelerometers always measure acceleration with respect to a reference frame of calibration that must be assumed to be inertial in the first place.
@@
Leibniz’s relative space, time and motion ideas were further developed by Ernst Mach in his 1883 The Science of Mechanics, in which he criticized Newton’s conclusion of his bucket of water experiment stating that “No one is competent to say how the experiment would turn out if the sides of the vessel increased in thickness and mass till they were ultimately several leagues thick.[Mach1881]. According to Mach, the relative rotation of the water with respect to the bucket produces no noticeable centrifugal forces, and such forces are instead produced by its relative rotation with respect to Earth and the other celestial bodies (in this way, the relative motion between the several leagues thick bucket and the water could produce noticeable centrifugal forces in the water, and perhaps their non-relative motion could reduce these forces when both rotate with respect to the rest of the universe). In contrast to Newton, who attempted to explain the physical effects of inertia through a sort of resistance to motion within an unobservable absolute space with no physical properties, Mach conceived inertia as an interaction that required other external bodies to manifest. Inertia would depend upon an interaction between masses. Newton did not suspect that the change in the water’s shape could be due to rotation relative to the rest of the universe because his action at a distance gravitational force from an infinite and homogeneous universe acting on the water cancels out according to Newton’s shell theorem and is independent of velocity or acceleration. But no local phenomena can ever be isolated from the rest of the universe, and its effect certainly reaches the water even if the sum of these forces is zero. According to Mach, and in opposition to Newton’s conclusions, a spinning bucket of water in an empty universe would not change the water’s shape (one could not detect relative motion in an empty universe, it would be undefinable), and if the rest of the universe was set spinning while the bucket was at rest, the water surface would curve, as both the spinning of the bucket and water or the spinning of the universe are indistinguishable systems without an absolute space. Thus, for Mach, the relativity symmetry of acceleration is broken by the existence of a background matter universe.
@@
Moreover, there is an exact coincidence between the local measurement of the angular velocity of the Earth by Foucault’s pendulum and the cosmological measurement through the apparent movement of distant stars and galaxies, which Newtonian gravity and mechanics cannot explain because it does not causally connect both measurements (and consider the determination of inertial frames by the fixed stars and Foucault’s pendulum a coincidence), but constitutes the basic idea of Mach’s principle: “The universe, as represented by the average motion of distant galaxies, does not appear to rotate relative to local inertial frames.[Bondi1997]. For Mach, its the distant masses the ones which hold the pendulum fixed with respect to the Earth.
@@
According to Mach, and known as Mach’s principle, the inertia of a body is not an independent and intrinsic property of matter (unlike in Newtonian mechanics, where a particle in an empty universe has the usual inertial properties), but rather the result of the action of the universe as a whole. Mach suggested that the fixed background distribution of matter in the universe must exert the inertial forces on a local accelerating body. In this way, the reference system with respect to which the universe is at rest or in uniform and rectilinear motion is a true inertial reference system. In Ernst Mach words, “I have remained to the present day the only one who insists upon referring the law of inertia to the Earth, and in the case of motions of great spatial and temporal extent to the fixed stars.[Mach1881]. Hence, inertial frames should be defined with respect to this rest frame, and local physical laws must be determined by the large-scale structure of the universe.
@@
Einstein’s pursuit of a relativistic gravitational theory was inspired by Mach, and he envisioned GR to fulfill his interpretation of Mach’s principle: “the metric tensor is completely caused and determined by the stress-energy tensor[Einstein1918]. He noted that following Mach, it was expected that “1. The inertia of a body must increase when ponderable masses are piled up in its neighborhood, 2. A body must experience an accelerating force when neighboring masses are accelerated, and, in fact, the force must be in the same direction as the acceleration, 3. A rotating hollow body must generate inside of itself a “Coriolis field”, which deflects moving bodies in the sense of the rotation, and a radial centrifugal field as well[Einstein1922]. He had hopes that by using tensors in GR, he would achieve the strong Machian version of relativity (in which the gravitational influence of the whole universe gave rise to inertia). But although consequences 2. and 3. were satisfied in GR (at least phenomenologically by frame-dragging), the first condition was not, because the immediate consequence of it is that a body in an otherwise empty universe should have no inertia, and the empty solution where the stress-energy tensor is zero everywhere in GR is the flat Minkowski spacetime of SR, in which test bodies have the usual inertia. He also referred to this consequence in 1913 stating that “It must be demanded that the inertial resistance of a body could be increased by having unaccelerated inertial masses arranged in its vicinity; and this increase of the inertial resistance must disappear again if these masses accelerate along with the body”. GR can be considered more Machian, since the metric defining inertia is dynamical and can be acted upon (in contrast to the case of Newton’s or SR, where the metric cannot be affected, and thus, its an absolute). Einstein then introduced the cosmological constant term (as the mass density of the universe) in his field equations in the hope that, with the cosmological term, they would have no solutions for a zero stress-energy tensor. However, de Sitter found in 1917 a solution for the field equations with the cosmological constant and with zero stress-energy tensor for an expanding universe, Einstein dismissed the cosmological constant as no longer justified, and he abandoned the ideas of Mach. In his effort to render GR fully Machian, Einstein reached the conclusion that the metric should diverge at infinity (in particular, that the time component of the metric tensor, associated with the Newtonian gravitational potential, should should increase without bound at infinity, and the spatial component should vanish) to satisfy the relativity of inertia (Mach’s principle): “It seems, therefore, that such a degeneration of the coefficients [of the components of the metric tensor], is required by the postulate of relativity of all inertia. This requirement implies that the potential energy $m\sqrt{B}$ [with $ds^{2}=-Ad\vec{x}^{2}+Bdt^{2}$] becomes infinitely great at infinity.[Einstein1917]. Unfortunately, Einstein thought at the time that the stars at infinity were moving very slowly, making the divergence of the metric incompatible with this belief, and dismissed the divergent metric for the cosmological constant due to his lack of knowledge of the expansion of the universe.
@@
Einstein’s equivalence principle (a frame linearly accelerated relative to an inertial frame in the Minkowskian spacetime of SR is locally identical to a frame at rest in a uniform gravitational field) axiomatically assumes that inertial and gravitational masses are equal. Gravity is the only infinite range force that cannot be screened due to the existence of only positive masses, unlike in electromagnetism, where positive and negative charges result in balance and neutrality, and astrophysical and cosmological structures are therefore not governed by electromagnetism. Moreover, there always exists a reference frame in which inertial forces vanish, just as for the case of gravitational forces in free fall, and the gravitational field, unlike the electromagnetic field, carries the information about a particle’s rest mass. These equivalences indicate that both inertial and gravitational forces are of the same nature. Following Mach, in any two-body interaction, the influence of all other matter inside their causal spacetime should be taken into account. Inertia is thus a form of gravitational induction, appearing when a body is accelerated with respect to the rest of the universe. But Newtonian gravity and GR were developed without considering today’s knowledge of the mass content and size of the observable universe. Then, it seems that to describe, for example, the Earth’s motion around the Sun, only local masses and distances are required, but the Machian perspective implies that the universe’s action is already taken into account in the Newtonian laws. The only way this can be true is through the Newtonian gravitational constant, which, together with the postulate of the inertial frame of absolute space, both constitute the two arbitrary choices of Newton.
@@
In Newtonian gravity, the gravitational constant $G$ is the only fundamental constant of the theory, and it is known with far less precision than any other fundamental dimensional constant in physics. The uncertainty in its measurement, together with its problematic in Quantum Field Theory with the Planck energy (which is used as a cutoff for the energy density of the vacuum and the theoretical quantum corrections estimation of the Higg’s mass, leading to the vacuum catastrophe and hierarchy problems) has led to questioning its constant nature repeatedly over the last century; see [Uzan2011] for constraints. GR carries this constant (a varying $G$ would imply a violation of the strong equivalence principle, but only the weak principle is supported by the very precise Eötvös experiments) with the addition of the speed of light, which can be derived from the vacuum permittivity and permeability constants. Therefore, it is reasonable to question whether the gravitational constant is also a derived parameter.
@@
The first historically suggestion of gravity arising necessarily as a consequence of the relativity of inertia came from Hans Reissner [Reissner1915]. Schrödinger soon identified in 1925 [Schrodinger1925] that GR did not fully implement Mach’s principle, and proposed a relationship between the gravitational potential of distant masses and the speed of light: “This remarkable relationship states that the (negative) potential of all masses at the point of observation, calculated with the gravitational constant valid at that observation point, must be equal to half the square of the speed of light.”. This conclusion was reached by imposing that the kinetic energy term had an origin in a potential-like interaction (such as all other forms of energy, which have an origin in an interaction) and was relational, so that $K=mv^{2}/2$ can be expressed through the gravitational Newtonian potential $Gm/r$ so that $K=(Gm_{i}/r_{ij})(m_{j}v_{ij}^{2}/2c^{2})$ with the need to introduce the square of the speed of light. The term $Gm_{i}/r_{ij}c^{2}=GM_{u}/c^{2}R_{u}\sim 1$ is considered missing in the original Newtonian formulation, with $M_{u}$ being the sum of all masses within $R_{u}$, the distance between the moving particle $m_{j}$ and $M_{u}$. Schrödinger’s formulation predicted a finite observable universe, since the potential of an infinite non-expanding eternal universe would also be infinite, even though the expansion of the universe was not known at the time. This relationship was later credited to Reissner’s 1915 work by Schrödinger himself.